# Universal and Experiment-calibrated Prediction of XANES through Crystal Graph Neural Network and Transfer Learning Strategy

### v1

@@ III Discussion
Different ways to achieve universal XANES prediction are compared in this work. Directly training one CGXAS_Uni model on a universal XANES dataset shows better performance for XANES prediction of almost all the elements than training a CGXAS_Spec_X model on an element-specific XANES subset. Since the element-specific subset used in CGXAS_Spec_X comes completely from the universal dataset for a specific element X, the enhanced XANES prediction performance of the CGXAS_Uni model must be attributed to other subsets of non-X elements in the universal dataset. Although the XANES spectra of different elements are significantly different regarding their energy ranges, the underlying physical rules for X-ray absorption and electron excitation are similar. Training model with the universal XANES dataset enables the neural network to learn the general underlying structure-XANES relationship regardless of the energy range for each element, thus enhancing the overall prediction accuracy for all elements. The fine-tuned CGXAS_FT_X model on the element-specific dataset has little influence on the prediction accuracy, implying that the knowledge of each element has already been well learned in the CGXAS_Uni model. In the previous work for 3d transition metal XANES prediction, Kharel et al. came out with an opposite result: the element specific model outperforms the universal model[RN365]. The cause the opposite results can be understood by the different structure features representations adopted in these two works. In this work, a GNN encoder is trained to process the structure information in the crystal graph, where a universal dataset containing more diverse structures is helpful to achieve better structure representations. In Kharel’s work, the universal dataset has no advantage in achieving better structure representations since a pretrained model is used for structure features encoding. On the contrary, the XANES prediction in universal model is more complex than that in element specific model, therefore yielding an opposite result. But pretraining on universal dataset can still reduce the overfitting risk for the finetuned universal model, therefore their finetuned universal models has better performance than the element specific models in general.
@@
It should be noted that the three experiment-calibrated models are developed for the purpose of proof-of-principles in this work. There some still some problems need to be solved before they can be put into practical use. For the peak width and some finer peaks in the high energy region, the experiment-calibrated models may give relatively large errors in the prediction. And there are also a few spectra with large discrepancy with experimental spectra. In future study, the universal, efficient, and experiment-based prediction of XANES can be further enhanced by obtaining more high-fidelity XANES data and developing more powerful AI algorithms, enabling the scientist to gain deeper insights into more complex scientific problems using XAS techniques.
@@ IV.1 Node and edge features representation
In the crystal graph, one atom is connected with the neighboring atoms within the cutoff radius of 8 Å through a directional edge. And the maximum number of neighbors per atom is 12 in this work. To keep the rotational invariance and encode the geometry information, the Gaussian basis set and the plane wave basis set are adopted as the edge features[RN158].
@@ IV.2 Model architecture of CGXAS
The model architecture of CGXAS is shown in Fig.6(b). The node features of the input graph are firstly embedded by an embedding layer. Then, the information of neighboring nodes is gathered to each node through the edge in the following graph convolution modules. Here, one graph convolution module is composed of a gated convolution layer of geo-CGNN[RN158], a normalization layer and a one layer perceptron. The updated node features of each graph convolution module $\omega^{i}$ are multiplied by $\frac{1}{i!}$ and summed over as the site vectors. The above modules together compose the GNN module in Fig.1. Next, the site vectors are processed by the multilayer perceptron and a Relu layer to generate the intensity of XANES of each absorption site. The activation function of the multilayer perceptron is Silu in this work. And the Relu layer is designed to avoid the prediction of negative intensity. Finally, the mask pooling layer will average the XANES over the unmasked sites, which is used to obtain the XANES of target site or specific element. The depths of gated convolution layers and multilayers perceptron are both five. The widths of input node features, hidden layer features and output spectrum are 146, 256 and 100 separately.
Figure 6: Crystal graph representation and detailed architecture of CGXAS. (a) Node features of crystal graph. (b) Model architecture of CGXAS.
@@ IV.3 XANES dataset
The dataset of simulated K edge XANES spectra used in this work covers the elements from the second to fifth period except the rare gas elements. The crystal structures of materials are derived from the Materials Project[RN408, RN409] using the pymatgen[RN289] and the XANES are obtained from an opensource XAS dataset generated through FEFF9[RN219]. Each XANES spectrum is matched with a distinct atom site in a material and is indexed by a unique id in the following format: “mp id - site index - spectrum type - edge type - absorption element”. For example, “mp-561304-56-XANES-K-S” represents the S K edge XANES of the 56th site of material “mp-561304” in Materials Project. For all the XANES spectra of the same element, a unified absorption energy range that can cover the pre-edge and near edge regions is chosen. And the XANES intensities at 100 uniform energy grid points in the corresponding element energy range are defined as the output of CGXAS. By fixing the absorption energy range of different elements, the multiple elements XANES can be predicted in the same neural network.
The experimental data used in transfer learning contains S K edge XANES of 48 different materials, Ti K edge XANES of 40 different materials, and Fe K edge XANES of 45 different materials. The spectra are obtained from beamline BL16U1 of the Shanghai Synchrotron Radiation Facility (SSRF), beamline 4B7A of Beijing Synchrotron Radiation Facility (BSRF), the XASLIB database (https://xaslib.xrayabsorption.org/), ESRF ID21 Sulfur XANES spectra database(https://www.esrf.fr/home/UsersAndScience/Experiments/XNP/ID21/php.html), MDR XAFS DB[RN404], and several literatures[RN122, RN293, RN294, RN295, RN297, RN298, RN350, RN331, RN360, RN361, RN363, RN357, RN396, RN321, RN402, RN397, RN400, RN398, RN401, RN399].
@@ IV.5 Transfer learning using experimental XANES
In the transfer learning of CGXAS_Exp_X, the pretrained CGXAS_Uni on massive simulated XANES is fine-tuned on the experimental XANES. To avoid over fitting, only the parameters of MLP are optimized layer by layer while the other parts of network remain the same as the pretrained model. The loss metric used in transfer learning is RSE. And the size of training, validation and testing set have a ratio of 6:1:1. The Adam optimizer and an initial learning rate of 0.0001 are chosen to optimize the layer of MLP. The maximum number of training epochs of each layer is 30.

### v2

@@ III Discussion
Different ways to achieve universal XANES prediction are compared in this work. Directly training one CGXAS_Uni model on a universal XANES dataset shows better performance for XANES prediction of almost all the elements than training a CGXAS_Spec_X model on an element-specific XANES subset. Since the element-specific subset used in CGXAS_Spec_X comes completely from the universal dataset for a specific element X, the enhanced XANES prediction performance of the CGXAS_Uni model must be attributed to other subsets of non-X elements in the universal dataset. Although the XANES spectra of different elements are significantly different regarding their energy ranges, the underlying physical rules for X-ray absorption and electron excitation are similar. Training model with the universal XANES dataset enables the neural network to learn the general underlying structure-XANES relationship regardless of the energy range for each element, thus enhancing the overall prediction accuracy for all elements. The fine-tuned CGXAS_FT_X model on the element-specific dataset has little influence on the prediction accuracy, implying that the knowledge of each element has already been well learned in the CGXAS_Uni model. In the previous work for 3d transition metal XANES prediction, Kharel et al. came out with an opposite result: the element specific model outperforms the universal model[25]. The cause the opposite results can be understood by the different structure features representations adopted in these two works. In this work, a GNN encoder is trained to process the structure information in the crystal graph, where a universal dataset containing more diverse structures is helpful to achieve better structure representations. In Kharel’s work, the universal dataset has no advantage in achieving better structure representations since a pretrained model is used for structure features encoding. On the contrary, the XANES prediction in universal model is more complex than that in element specific model, therefore yielding an opposite result. But pretraining on universal dataset can still reduce the overfitting risk for the finetuned universal model, therefore their finetuned universal models has better performance than the element specific models in general.
@@
It should be noted that the three experiment-calibrated models are developed for the purpose of proof-of-principles in this work. There some still some problems need to be solved before they can be put into practical use. For the pre-edge structure and some finer peaks in the high energy region, the experiment-calibrated models may give relatively large errors in the prediction. And there are also a few spectra with large discrepancy with experimental spectra. In future study, the universal, efficient, and experiment-based prediction of XANES can be further enhanced by obtaining more high-fidelity XANES data and developing more powerful AI algorithms, enabling the scientist to gain deeper insights into more complex scientific problems using XAS techniques.
@@ IV.1 Node and edge features representation
In the crystal graph, one atom is connected with the neighboring atoms within the cutoff radius of 8 Å through a directional edge. And the maximum number of neighbors per atom is 12 in this work. To keep the rotational invariance and encode the geometry information, the Gaussian basis set and the plane wave basis set are adopted as the edge features[9].
@@ IV.2 Model architecture of CGXAS
The model architecture of CGXAS is shown in Fig.6(b). The node features of the input graph are firstly embedded by an embedding layer. Then, the information of neighboring nodes is gathered to each node through the edge in the following graph convolution modules. Here, one graph convolution module is composed of a gated convolution layer of geo-CGNN[9], a normalization layer and a one layer perceptron. The updated node features of each graph convolution module $\omega^{i}$ are multiplied by $\frac{1}{i!}$ and summed over as the site vectors. The above modules together compose the GNN module in Fig.1. Next, the site vectors are processed by the multilayer perceptron and a Relu layer to generate the intensity of XANES of each absorption site. The activation function of the multilayer perceptron is Silu in this work. And the Relu layer is designed to avoid the prediction of negative intensity. Finally, the mask pooling layer will average the XANES over the unmasked sites, which is used to obtain the XANES of target site or specific element. The depths of gated convolution layers and multilayers perceptron are both five. The widths of input node features, hidden layer features and output spectrum are 146, 256 and 100 separately.
Figure 6: Crystal graph representation and detailed architecture of CGXAS. (a) Node features of crystal graph. (b) Model architecture of CGXAS. (c) Flow chart of transfer learning.
@@ IV.3 XANES dataset
The dataset of simulated K edge XANES spectra used in this work covers the elements from the second to fifth period except the rare gas elements. The crystal structures of materials are derived from the Materials Project[18, 20] using the pymatgen[36] and the XANES are obtained from an opensource XAS dataset generated through FEFF9[33]. Each XANES spectrum is matched with a distinct atom site in a material and is indexed by a unique id in the following format: “mp id - site index - spectrum type - edge type - absorption element”. For example, “mp-561304-56-XANES-K-S” represents the S K edge XANES of the 56th site of material “mp-561304” in Materials Project. For all the XANES spectra of the same element, a unified absorption energy range that can cover the pre-edge and near edge regions is chosen. And the XANES intensities at 100 uniform energy grid points in the corresponding element energy range are defined as the output of CGXAS. By fixing the absorption energy range of different elements, the multiple elements XANES can be predicted in the same neural network.
The experimental data used in transfer learning contains S K edge XANES of 48 different materials, Ti K edge XANES of 40 different materials, and Fe K edge XANES of 45 different materials. The spectra are obtained from beamline BL16U1 of the Shanghai Synchrotron Radiation Facility (SSRF), beamline 4B7A of Beijing Synchrotron Radiation Facility (BSRF), the XASLIB database (https://xaslib.xrayabsorption.org/), ESRF ID21 Sulfur XANES spectra database(https://www.esrf.fr/home/UsersAndScience/Experiments/XNP/ID21/php.html), MDR XAFS DB[19], and several literatures[17, 3, 26, 10, 55, 32, 13, 15, 12, 14, 45, 27, 2, 52, 4, 11, 30, 31, 41, 56].
@@ IV.5 Transfer learning using experimental XANES
The transfer learning process (shown in Fig.6(c)) can be divided into the following step: 1. The experimental dataset is firstly used to finetune the CGXAS_Uni model to a initial transferred model CGXAS_Exp_X. 2. The CGXAS_Exp_X model is applied to calibrate the energy of simulated dataset. 3. The calibrated simulated dataset is then used to finetune the CGXAS_Uni model to the CGXAS_Cali_X model. 4. CGXAS_Cali_X model is further finetuned on the experimental dataset to the new CGXAS_Exp_X model. 5. Repeat step 2-4 for 10 times and select the CGXAS_Exp_X model with the lowest edge energy error in testing set. To avoid over fitting, only the parameters of MLP are optimized layer by layer while the other parts of network remain the same as the pretrained model. The loss metric used in transfer learning is RSE. And the size of training, validation and testing set have a ratio of 6:1:1. The Adam optimizer and an initial learning rate of 0.0001 are chosen to optimize the layer of MLP. The maximum number of training epochs of each layer is 30.
The similarity of testing set compared with training and validation is measured by the distance of testing sample to the training and validation set (DTV):
$\text{DTV}=\min\{\sqrt{(f_{1,i}-f_{1,t})^{2}+(f_{2,i}-f_{2,t})^{2}}\}$ (3)
where $f_{1,i}$ and $f_{2,i}$ are the two t-SNE features of the $i^{\text{th}}$ sample in training and validation set, n is the number of samples in training and validation set, and $f_{1,t}$ and $f_{2,t}$ are the two t-SNE features of the sample in testing set.